OpenStudy (anonymous):

Prove cot x cos = cos^2 x csc

OpenStudy (john_es):

\[\cot x\cos x=\frac{\cos x}{\sin x}\cos x=\cos^2x\frac{1}{\sin x}=\cos^2x\csc x\]